OpenStudy (anonymous):

What is the simplified form of the sixth root of x times the sixth root of x times the sixth root of x times the sixth root of x? x to the two-thirds power x to the three halves power 4 times the sixth root of x the sixth root of 4x

OpenStudy (anonymous):

\[\sqrt[6]{x^4}=\sqrt[3]{x^2}\]

OpenStudy (anonymous):

or \[x^{\frac{2}{3}}\]

OpenStudy (anonymous):

What is the remainder when (3x4 + 2x3 - x2 + 2x - 24) ÷ (x + 2)? 0 5 10 15

OpenStudy (anonymous):

\[f(-2)=0\] so 0

OpenStudy (anonymous):

A function is created to represent the miles per gallon your car gets. What restrictions would be made to the domain? The domain would only include integers. The domain would only include positive integers. The domain would include all real numbers. The domain would only include positive numbers.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

positive numbers of course

OpenStudy (anonymous):

which seems to be two choices

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

no D

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

What is the equation of the quadratic graph with a focus of (3, 6) and a directrix of y = 4? f(x) = one fourth (x − 3)2 + 1 f(x) = one fourth (x − 3)2 + 5 f(x) = −one fourth (x − 2)2 + 5 f(x) = −one fourth (x − 2)2

OpenStudy (anonymous):

Ella runs a cat rescue center. She started with 2 cats, and her center has an increase of 5 cats per year. Which of the graphs below represents the number of cats per year at Ella's center? graph of line going through 2, 5 and 5, 8 graph of line going through 1, 7 and 2, 12 graph of line going through 1, 1 and 3, 3 graph of line going through 1, 3 and 2, 8

OpenStudy (anonymous):

has to go through 1 7 since in the first year they get 5 cats and already have two

OpenStudy (anonymous):

Find the solution of the square root of the quantity of x plus 3 plus 4 equals 6 and determine if it is an extraneous solution. x = 1; not extraneous x = 1; extraneous x = 29; extraneous x = 29; not extraneous

OpenStudy (anonymous):

\[\sqrt{x+3}+4=6\]?

OpenStudy (anonymous):

thne \[\sqrt{x+3}=2\\ x+3=4\\ x=1\] is the "not extraneous" solution

OpenStudy (anonymous):

yep u got it

OpenStudy (anonymous):

lol imagine

OpenStudy (anonymous):

Two mechanics, Martin and Gordon, are working on your car. Martin can complete the work in 4 hours, while Gordon can complete the work in 2 hours. How many hours does the maintenance take if they work together? one half three fourths six eighths four thirds

OpenStudy (anonymous):

\[\frac{4\times 2}{4+2}\]

OpenStudy (anonymous):

\[\frac{8}{6}=\frac{4}{3}\]

OpenStudy (anonymous):

Given the function f(x) = The quantity of 2x minus 5, divided by 3, which of the below expressions is correct? f−1(x) = The quantity of 5 minus 2x, divided by 3. f−1(x) = The quantity of 3x minus 5, divided by 2. f−1(x) = The quantity of 3x plus 5, divided by 2. f−1(x) = The quantity of negative 2x minus 5, divided by 3.

OpenStudy (anonymous):

igot the question marks gotta refresh

OpenStudy (anonymous):

\[f(x)=\frac{2x-5}{3}\]?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

What is the simplified form of x minus 2 over x squared plus x minus 6 ÷ x squared plus 5x plus 4 over x plus 4 ? 1 over the quantity x plus 3 times the quantity x plus 4 1 over the quantity x plus 3 times the quantity x minus 2 1 over the quantity x plus 4 times the quantity x minus 2 1 over the quantity x plus 3 times the quantity x plus 1

OpenStudy (anonymous):

question marks again

OpenStudy (anonymous):

x minus 2 over x squared plus x minus 6 ÷ x squared plus 5x plus 4 over x plus 4 ?

OpenStudy (anonymous):

\[\frac{x-2}{x^2}+\frac{x-6}{x^2}+\frac{5x+4}{x+4}\]?

OpenStudy (anonymous):

is that it?

OpenStudy (anonymous):

no i'll draw it

OpenStudy (anonymous):

can you post a screen shot? is i really much quicker

OpenStudy (anonymous):

i take it it is on line

OpenStudy (anonymous):